# Quantum-centric Supercomputing for Materials Science: A Perspective on Challenges and Future Directions

### v1

@@ IV.1 Why use classical simulators
IV.2 Quantum circuits simulators
@@
Before we start, we note that not every quantum circuit is difficult to simulate on a classical computer. Trivially, a quantum circuit that carries out a purely classical computation is not difficult to simulate on a classical computer. Indeed, some of the quantum circuits that are not difficult to simulate include Clifford circuits [356, 357, 358] and the quantum Fourier transform [359, 360], to name a couple. Some quantum circuits are known to be classically simulable for specific input states as well [358]. If the simulation does not need to produce the full spectrum of the output state of a given quantum computation but rather a sampling of such, as it would be the case for obtaining a final, classical result out of a quantum computer upon measurements, the difficulty of the classical simulation can dramatically decrease [361]. Further decrease in the difficulty is obtained if the simulation is to mimic an error-prone quantum computer executing a quantum circuit [362]. We further note in passing that simulation of a quantum circuit that exhibits sparse couplings between densely connected sets of qubits can be more amenable to classical simulations [363], using methodologies not unlike circuit cutting discussed in Sec. II.6.1.
IV.2.1 State-vector simulators
@@
IV.2.2 Density matrix simulators
@@
IV.2.3 Tensor network simulators
@@
There are different types of tensor network simulators: Matrix Product States (MPS), Projected Entangled Pair States (PEPS), Tree Tensor Networks (TTN), Multi-scale Entanglement Renormalization Ansatz (MERA), and Continuous Variable Tensor Network (CVTN) to name a few as shown on Figure 8. MPS is arguably one of the most popular simulators, and it is particularly useful for simulating one-dimensional quantum systems. In MPS, the tensors are arranged in a chain, with each tensor representing the state of a qubit and the connections between them capturing the nearest-neighbor entanglement. One of the MPS applications is an efficient tensor network simulation of lossy Gaussian boson sampling experiments, which has been recently demonstrated [380, 381].
@@
  • CTVN simulators are tailored for quantum systems with continuous variables, like quantum fields or modes of light. They are adept at handling systems where particle number isn’t conserved and are crucial in studying non-Gaussian states and processes in quantum optics and field theories.
@@
IV.2.4 Open system Lindblad quantum simulators
@@
IV.3 Overview of classical simulators
@@
V Use case identification
V.1 Criteria for use case identification
@@ V.2 Overview of use cases
V.3 Relevant use cases
V.3.1 Ground-state electronic structure
@@
[…]
@@ V.3.2 Embedding
  • Density-based embedding theories define an active space in real space and partition the density of the system into the active space and the environment. The latter is described at the DFT level, and the former is described either using DFT with a higher-level exchange-correlation functional (DFT-in-DFT) or with wavefunction-based quantum chemistry methods (WF-in-DFT).
@@
  • Embedding theories based on Green’s function define an active space by a set of single-particle electronic states. The non-local, frequency-dependent self-energy of the active space is expressed as a sum of terms evaluated at high and low levels of theory, with an additional term to remove the double counting. Examples of such theories include dynamical mean-field theory embedding (DMFT+DFT [431, 432] or DMFT+GW [433, 434, 435, 436, 437, 438]), self-energy embedding theory (SEET) [439, 440, 441], and quantum defect embedding theory (QDET) [442, 443, 444, 445]. They mainly differ in the choice of the high-level and low-level theories and the partitioning of the total self-energy into the active space and the environment.
[…]
V.3.3 Model systems
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
On classical computers, the Emery model has been investigated with several numerical methods, including direct simulations of lattices by exact diagonalization of up to 24 sites [496], quantum Monte Carlo and density matrix renormalization group (DMRG) of 80 and 144 sites [497, 498], and embedding methods like density matrix embedding theory with 12 sites [499]. These simulations have been primarily directed at simulating ground-state spin-spin, and $d$-wave correlation functions, to establish e.g. the presence of stripe or $d$-wave superconducting order [496, 497, 498, 499].
[…]
V.4 Excited states
@@
V.5 Applications of electronic structure
V.5.1 Vibrational structure calculations
@@
V.5.2 Coarse-grained simulations
@@
V.5.3 An example use case: quantum metamaterials
@@ VI Conclusions and Outlook
The different sections have highlighted the challenges at various levels. The first two sections closely tie together, listing fundamental algorithms for simulation of quantum systems and specific hurdles that arise in implementing them at scale both for noisy quantum devices and fault-tolerant architectures, with an emphasis on the classical workloads. Then, we moved our focus closer to the architecture design arguing in details about the difficulties that are faced when optimizing quantum and classical workloads coming from hybrid algorithms. Before tying up all the threads in the section about use cases, we have discussed how classical simulations of quantum systems can be used both for approximate verification and to identify hard use cases. Finally, we have provided criteria for the identification of good use cases in materials science, drawing upon the previous sections. Based on these criteria, we have discussed a few exemplary cases, representative of the variety of applications of materials science in mind.
One major takeaway from this perspective is to suggest how we can lay the grounds to think about quantum computing to work in synergy with classical high-performance computing in quantum-centric supercomputing centers. Materials science provides a great setting for use cases, which have potential for quantum advantage for scientific and industrial applications.

### v2

@@ VI.1 Why use classical simulators
VI.2 Quantum circuits simulators
@@
Before we start, we note that not every quantum circuit is difficult to simulate on a classical computer. Trivially, a quantum circuit that carries out a purely classical computation is not difficult to simulate on a classical computer. Indeed, some of the quantum circuits that are not difficult to simulate include Clifford circuits [387, 388, 389] and the quantum Fourier transform [390, 391], to name a couple. Some quantum circuits are known to be classically simulable for specific input states as well [389]. If the simulation does not need to produce the full spectrum of the output state of a given quantum computation but rather a sampling of such, as it would be the case for obtaining a final, classical result out of a quantum computer upon measurements, the difficulty of the classical simulation can dramatically decrease [392]. Further decrease in the difficulty is obtained if the simulation is to mimic an error-prone quantum computer executing a quantum circuit [393]. We further note in passing that simulation of a quantum circuit that exhibits sparse couplings between densely connected sets of qubits can be more amenable to classical simulations [394], using methodologies not unlike circuit cutting discussed in Sec. III.6.1.
VI.2.1 State-vector simulators
@@
VI.2.2 Density matrix simulators
@@
VI.2.3 Tensor network simulators
@@
There are different types of tensor network simulators: Matrix Product States (MPS), Projected Entangled Pair (PEPS), Tree Tensor Networks (TTN), Multi-scale Entanglement Renormalization Ansatz (MERA), and Continuous Variable Tensor Network (CVTN) to name a few as shown on Figure 9. MPS is arguably one of the most popular simulators, and it is particularly useful for simulating one-dimensional quantum systems. In MPS, the tensors are arranged in a chain, with each tensor representing the state of a qubit and the connections between them capturing the nearest-neighbor entanglement. One of the MPS applications is an efficient tensor network simulation of lossy Gaussian boson sampling experiments, which has been recently demonstrated [411, 412].
@@
  • CVTN simulators are tailored for quantum systems with continuous variables, like quantum fields or modes of light. They are adept at handling systems where particle number isn’t conserved and are crucial in studying non-Gaussian states and processes in quantum optics and field theories.
@@
VI.2.4 Open system Lindblad quantum simulators
@@
VI.3 Overview of classical simulators
@@
VII Use case identification
VII.1 Criteria for use case identification
@@ VII.2 Overview of use cases
VII.3 Relevant use cases
VII.3.1 Ground-state electronic structure
@@
    • the iron-molybdenum cofactor (FeMo-cofactor) of nitrogenase, which catalyzes the 6-electron reduction of N2 to ammonium in biological nitrogen fixation [431, 416], and can be modeled by a (113e,76o) active space [429, 432]
@@ VII.3.2 Embedding
  • Density-based embedding theories define an active space in real space and partition the density of the system into the active space and the environment. The latter is described at the DFT level, and the former is described either using DFT with a higher-level exchange-correlation functional (DFT-in-DFT) or with wavefunction-based quantum chemistry methods (WF-in-DFT) [452, 453, 454, 455].
@@
  • Embedding theories based on Green’s function define an active space by a set of single-particle electronic states. The non-local, frequency-dependent self-energy of the active space is expressed as a sum of terms evaluated at high and low levels of theory, with an additional term to remove the double counting. Examples of such theories include dynamical mean-field theory embedding (DMFT) (DMFT+DFT [466, 467] or DMFT+GW [468, 469, 470, 471, 472, 473]), self-energy embedding theory (SEET) [474, 475, 476], and quantum defect embedding theory (QDET) [477, 478, 479, 480]. They mainly differ in the choice of the high-level and low-level theories and the partitioning of the total self-energy into the active space and the environment.
In practice, a quantum embedding calculation typically starts with a low-level calculation of the full system, the results of which are subsequently employed to determine parameters of the many-body effective Hamiltonian of the active space. In QDET, for example, the effective Hamiltonian comprises a one-body term and a two-body term. In the one-body term, an exact double counting correction has been derived for the case where G0W0 is the low-level theory. In the two-body term, the effect of the environment is incorporated through an effective screening. The QDET Hamiltonian can be computed without any explicit summation over empty states, enabling its application to large systems. Leveraging GPU-accelerated classical supercomputers, it is feasible to simulate systems consisting of over a thousand atoms and active spaces consisting of over a hundred orbitals. The effective Hamiltonian, written in second quantization, can be diagonalized on classical computers using FCI, or on quantum computers [478, 481, 482, 483, 484, 485, 486, 487, 248] by mapping it to qubits and quantum gates and finding its lowest eigenstates using a suitable quantum eigensolver. Challenges remain in facilitating the communication of data from classical (super)computers to quantum computers, accommodating the effective Hamiltonian with the available number of gates and circuit depth, and mitigating noises present on near-term computers which may lead to unphysical results. Intriguingly, quantum computers themselves are employed to investigate the properties of spin defects in solids. These defects hold promise as prospective candidates for the implementation of improved quantum computers, thus forming a positive feedback loop at the intersection of computational materials science and quantum computing.
Additionally, CAS-DMET and NEVPT2-DMET described before have been effectively used to examine the adsorption and emission spectra of the negatively charged nitrogen-vacancy (NV-1) and the neutral silicon vacancy (SiV0) defects in diamond, as well as the neutral oxygen vacancy (OV0) defect in both bulk and surface magnesium oxide, as reported in recent literature. [201, 202, 203] In particular, for the NV-1 defect in diamond, NEVPT2-DMET with a significantly reduced embedding space and a complete active space (CAS) configuration of 10 electrons in 9 orbitals (denoted as 10e, 9o) accurately predicts a triplet-triplet excitation energy of 2.31 eV. This is in close agreement with the experimentally observed value of 2.18 eV. Additionally, it predicts an optically inactive singlet-singlet transition at 1.02 eV, closely approximating the experimental value of 1.26 eV. For the neutral OV0 defect in MgO bulk, NEVPT2-DMET, with a CAS configuration of 2e, 5o, estimates the excitation energy to be 5.24 eV, aligning well with the experimental absorption maximum of 5.03 eV. Studies of the optically allowed $S_{0}\rightarrow S_{1}$ transition for the OV0 defect on the MgO surface have also been conducted using CAS-DMET, NEVPT2-DMET, and DME-PDFT methods. However, varying experimental results ranging from 1-5 eV have been reported, which makes direct comparison difficult. Using MP2 and CCSD solvers for the DMET embedding subspace, the adsorption energies for CO on MgO have been calculated using the largest embedding subspaces, and these have been found to be within 1.2 kcal/mol of the non-embedding reference values. [230] So far, the LASUCC and LASSI methods have primarily been tested on strongly correlated molecular systems and has been utilized to study the dissociation of (H2)2 into two H2 molecules, the dissociation of the two double bonds in trans-butadiene, and the $J$-coupling in tris-($\mu$-hydroxo)-bridged chromium compound model [233, 234]. The accuracy of these studies has been confirmed through excellent agreement with the results from corresponding CASCI calculations performed using LAS orbitals.
VII.3.3 Model systems
@@
The dominant effective model for quantum particles in materials is band structure, and for metals, Fermi liquid theory. However, a major challenge is how this paradigm should be altered when it is no longer a good description of the physical system. Examples of these include the high-Tc cuprates and other transition-metal compounds like spin liquids, which do not appear to be well-described by these simple effective Hamiltonians. For these systems, many models have been proposed, such as the Emery and Heisenberg reviewed below. While these models have been extensively studied analytically and numerically, and have significantly enhanced our understanding of the physics of correlated electrons, their effectiveness for describing a real complex system of interest is often unclear. At the same time, more complex effective models can be commensurately more difficult to solve, so one would like to also find an accurate effective model that is computationally tractable. Very importantly, most such models feature free parameters that need to be given precise values to be of relevance to the physics of high-Tc cuprates and spin liquids. These parameters are often assigned starting from experimental data and/or classical electronic structure simulations, leading to both biases and uncertainties.
@@
Figure 12: Left: crystal structure of $\alpha$-RuCl3, illustrating the van der Walls gap between the RuCl6 honeycomb layers (with Ru/Cl atoms represented as gray/green circles) and magnetic Ru3+ ions aligned as in the zigzag phase (red arrows). Right: schematic representation of the two-dimensional Heisenberg-Kitaev model with spins represented as white circles and $x/y/z$ bonds as green/blue/red links respectively. Quantum simulations to date have focused on 1 and 2 hexagons, corresponding to 6 and 10 spins respectively [502].
A concrete example of a material of interest as a proximate spin liquid [498] is ruthenium trichloride, $\alpha$-RuCl3 (see Fig. 12). This transition metal halide has a crystal structure made up of stacked honeycomb layers of edge-sharing RuCl6 octahedra. Plumb et al [503] found that spin-orbit coupling in this material is substantial, leading to a $j=1/2$ description of the Ru3+ valence electrons. Since this material is built up with edge-sharing RuCl6 octahedra, it is believed [504, 505] that its low-energy excitations are parametrized by the Kitaev-Heisenberg Hamiltonian
@@
where $\langle ij\rangle_{\gamma}$ denotes summation over the bonds $\gamma\in\{x,y,z\}$. This feature makes $\alpha$-RuCl3 a material of great interest in the ongoing search for a Kitaev spin-liquid ground state [506, 507, 508]. Although $\alpha$-RuCl3 orders magnetically at low temperature with zigzag magnetic order [508, 509], this material has shown some signatures of spin-liquid physics, such as a broad continuum of magnetic excitations identified in both Raman scattering [510] and inelastic neutron scattering measurements [508]. Therefore, the degree of similarity to the Kitaev model, as well as the interpretation of spectroscopic and thermal measurements in terms of modifications to the Kitaev model, is currently intensely debated. We remark that, in the literature [511], the Kitaev-Heisenberg Hamiltonian is augmented by additional couplings,
@@
with $\{\alpha,\beta,\gamma\}=\{x,y,z\},\{z,x,y\},\{y,z,x\}$ for $\gamma=z,y,x$ respectively. Several parameter configurations have been proposed to describe the Hamiltonian $H$, ranging from first-principle methods [512, 511] to phenomenological analyses [508, 511]. Recent literature has focused on $J=-1.53$, $K=-24.4$, $\Gamma=5.25$, $\Gamma^{\prime}=-0.95$ [502]. The coupling that is believed to be the leading one is the (negative) Kitaev term $K<0$. The off-diagonal exchange term $\Gamma>0$ is also believed to be significant, while the ferromagnetic exchange $J$ is believed to be subleading [513]. However, it should be noted that the parameters of $H$ vary quite significantly between the studies [514, 511]. The exactly solvable Kitaev model corresponds to $J=\Gamma=\Gamma^{\prime}=0$, and it is argued that, in the parameter regime of $\alpha$-RuCl3, both the excitations and the heat capacity show echoes of the two kinds of Majorana fermions that exist at the solvable point [515, 516].
The combination of rich physical behavior and challenging ground- and thermal-state preparation makes quantum spin liquid models like the Kitaev-Heisenberg Hamiltonian approximating $\alpha$-RuCl3 interesting and challenging problems for classical and quantum simulations. The presence of $2$-local spin-$1/2$ Hamiltonians makes them compelling targets for simulations on quantum devices, focused on the calculation of specific heats and dynamical structure factors relevant to understanding the low-energy excitation spectrum of the Hamiltonian.
@@
One concrete example of a fermionic lattice model that can be connected to a technological application is the three-band Hubbard model in two dimensions, also known as the Emery model [530]. This model is generally believed to contain the essential properties of high-Tc cuprate superconductors that arise from the interplay between the copper $d_{x^{2}-y^{2}}$ and oxygen $p_{x,y}$ orbitals in the CuO2 layers (see Figure 13). It is defined, in the so-called “phase representation”, by the Hamiltonian
@@
On classical computers, the Emery model has been investigated with several numerical methods, including direct simulations of lattices by exact diagonalization of up to 24 sites [531], quantum Monte Carlo and DMRG of 80 and 144 sites [532, 533], and embedding methods like density matrix embedding theory with 12 sites [534]. These simulations have been primarily directed at simulating ground-state spin-spin, and $d$-wave correlation functions, to establish e.g. the presence of stripe or $d$-wave superconducting order [531, 532, 533, 534].
Due to the complexity of the model, unlike in the one-band case, a consensus on much of the physics has yet to be reached, providing an opportunity for high-accuracy quantum computing algorithms to elucidate the properties of high-Tc cuprate superconductors by simulating lattices of O(100) sites.
The Emery model is often simplified considering a one-band Hubbard model. While much of the physics of high-Tc cuprate superconductors, e.g. $d$-wave pairing, density waves, pseudogap phase, stripe order, is described by the one-band Hubbard model Hamiltonian, there are still important reasons to go beyond the one-band picture to study the original three-band model directly. For instance, (a) some important physics may be lost in the reduction to the one-band approximation (such as a role for the oxygen degrees of freedom in the pseudogap phase [535]), (b) near degeneracies of competing states seen in the one-band case [529] may be resolved with the additional degrees of freedom of the three-band model, and (c) the three-band model retains the atomic structure of the CuO2 layer and thus has a direct link to the structure of real materials as well as experimental measurements of orders at the atomic scale [536, 537, 538].
VII.4 Excited states
@@
VII.5 Applications of electronic structure
VII.5.1 Vibrational structure calculations
@@
VII.5.2 Coarse-grained simulations
@@
VII.5.3 An example use case: quantum metamaterials
@@ VIII Conclusions and Outlook
The different sections have highlighted the challenges at various levels. The first two sections closely tie together, listing fundamental algorithms for simulation of quantum systems and specific hurdles that arise in implementing them at scale both for noisy quantum devices and fault-tolerant architectures, with an emphasis on the classical workloads. Then, we moved our focus closer to the architecture design arguing in details about the difficulties that are faced when optimizing quantum and classical workloads coming from hybrid algorithms. We discussed more general topics around quantum-centric supercomputing, namely, workload management and programming models, and how extensions to existing classical computing frameworks can help to address quantum computing device requirements. Before tying up all the threads in the section about use cases, we have discussed how classical simulations of quantum systems can be used both for approximate verification and to identify hard use cases. Finally, we have provided criteria for the identification of good use cases in materials science, drawing upon the previous sections. Based on these criteria, we have discussed a few exemplary cases, representative of the variety of applications of materials science in mind.
One major takeaway from this perspective is to suggest how we can lay the grounds to think about quantum computing together with classical high-performance computing, in quantum-centric supercomputing centers. Materials science provides a great setting for use cases, which have potential for quantum advantage in scientific and industrial applications.
@@ Acknowledgements.
ASIC
application-specific integrated circuit
BK
Bravyi-Kitaev
CAR
canonical anticommutation relations
CASSCF
complete active space self-consistent field
CCR
canonical commutation relations
CNOT
Controlled NOT gate
CVTN
Continuous Variable Tensor Network
DAG
directed acyclic graph
DFT
density functional theory
DMET
density matrix embedding theory
DMFT
dynamical mean-field theory embedding
DMRG
density matrix renormalization group
DUA
Data Use Agreements
FCI
Full Configuration Interaction
FEM
finite element method
FPGA
field programmable gate array
GPU
graphics processing unit
HHL
Harrow-Hassidim-Lloyd
HPC
high performance computing
IC
informationally incomplete
ISA
Instruction Set Architecture
JW
Jordan-Wigner
LASSCF
localized active space self-consistent field
LCU
linear combination of unitaries
MC-PDFT
multiconfiguration pair-density functional theory
MERA
Multi-scale Entanglement Renormalization Ansatz
MIMD
multiple instruction, multiple data
MPI
Message Passing Interface
MPS
Matrix Product States
MRCI
multireference configuration interaction
OpenMP
Open Multi-Processing
PDE
partial differential equation
PEC
probabilistic error cancellation
PEPS
Projected Entangled Pair
PES
potential energy surface
POVM
Positive Operator Valued Measure
QCSC
Quantum-centric Supercomputing
QDC
Quantum Data Center
QDET
quantum defect embedding theory
QEC
quantum error correction
QEM
quantum error mitigation
QMA
Quantum Merlin Arthur
QPE
Quantum Phase Estimation
QPU
quantum processing unit
QRAM
quantum random access memory
QROM
Quantum Read-Only Memory
RDM
reduced density matrix
SDK
software development kits
SIMD
single instruction, multiple data
SIMT
single instruction, multiple threads
SISD
single instruction, single data
TTN
Tree Tensor Networks
UCC
Unitary Coupled Cluster
VQA
variational quantum algorithm
VQE
Variational Quantum Eigensolver
WFMS
workflow management system
WMS
workload management systems
ZNE
Zero-Noise Extrapolation